OpenStudy (exoticinspirit):

What is the "!" in math? like 5! is it used in algebra 2?

OpenStudy (anonymous):

Factorials are very simple things. They're just products, indicated by an exclamation mark. For instance, "four factorial" is written as "4!" and means 1×2×3×4 = 24. In general, n! ("enn factorial") means the product of all the whole numbers from 1 to n; that is, n! = 1×2×3×...×n.

OpenStudy (black_ninja123):

5! would be 5*4*3*2*1=120 I don't know about algebra 2 but it is used in algebra 1

OpenStudy (exoticinspirit):

Thank you both so much makes a lot of sense :D

OpenStudy (anonymous):

np!

OpenStudy (black_ninja123):

Your welcome